# Current address: ]College of William and Mary, Williamsburg, Virginia 23187, USA Current address: ]Thomas Jefferson National Accelerator Facility, Newport News, Virginia 23606, USA

### v3

@@ VII.6.2 Isospin Study of
Our results on the isovector and isoscalar combination of from EG4 are shown in Fig. 26. We also provide results based on EG4 data and E97-110 (3He) data [104]. The total experimental systematic uncertainties on the proton and deuteron/neutron(3He) are combined quadratically, as are the statistical uncertainties. The low- uncertainties are added linearly. The results shown in Fig. 26 agree reasonably well with those from the previous experiments EG1b and EG1b/E94-010 [43]. However, unlike for , results from the EG4-only and EG4/E97-110 approaches agree only at the higher points. On the other hand, if we combine linearly the total systematic uncertainties of the two experiments, the tension at low vanishes.
@@ VII.7 Discussions
As discussed in previous sections, the success of EFT in describing our results on the moments of the spin structure function remains limited. One reason for this limited success may be coming from the difficulty to fully account for the degree of freedom in the calculation. In fact, early EFT calculations [152, 79] did not include explicitly the excitation, which slows the convergence of the EFT perturbation series, or they included it phenomenologically [81, 83]. This was thought to not be important for some observables for which the contribution was expected to be suppressed, such as the longitudinal-transverse spin polarizability and the isovector quantities [151] or . It therefore came as a surprise that much of the early nucleon spin structure function data [92] disagreed with calculations [152, 79, 81, 83], noticeably and . This puzzle prompted refined EFT calculations [59, 60] and a new experimental program at JLab optimized to cover the EFT domain [21, 71, 163], including the EG4 experiment reported here.
@@
The latest EFT calculations by Bernard et al. [59] and Alarcón et al. [60] both included the but differ in their expansion method for the (pion)- corrections. In EFT, the general expansion parameter is , where GeV and GeV is the chiral symmetry breaking scale. To explicitly account for the , the nucleon- mass difference GeV must be included in the chiral expansion. Bernard et al. [59] treated as a small parameter of the same order as . Alarcón et al. [60] used as an intermediate scale so that is the expansion parameter for the - corrections. This is referred to as the “-expansion scheme”. Additionally, Alarcón et al. included a phenomenological form factor, viz not derived within EFT, that suppresses the -dependence of observables at large . This clearly extends the range of applicability of the Alarcón et al. prediction [60] to higher compared to that of Bernard et al. [59], albeit by introducing an extra scale factor.
@@
An important outcome that can be drawn from the results shown in this section is that EFT, although successful in many instances, remains challenged by results from dedicated polarized experiments at low . This includes not only EG4, but also other experiments that used different detectors, methods, and spin observables [71, 104, 163]. To address this problem, further improvements on EFT would be required, although it is very difficult to upgrade the predictions to the next order. Since EFT is the leading QCD-rooted approach to calculating the effects of the strong force at large distances, this difficulty poses a problem in our pursuit of a complete description of nature. Alternatively or complementarily, it would be helpful to have predictions from other non-perturbative approaches to QCD, such as lattice QCD, the Dyson-Schwinger Equations, or the AdS/CFT methods. Recent progress in calculating doubly-virtual Compton scattering (VVCS) amplitudes in lattice QCD now makes moments involved in sum rules within reach of this method, see, e.g., Refs. [164, 165] for recent development. At present, only calculations for spin-independent VVCS amplitudes are available; these calculations are difficult because they involve computing 4-point correlation functions, in contrast to usual Lattice QCD calculations that require less computing-extensive 2- and 3-point correlations functions. Only recently, advances in methods and computer power have allowed such a computation. There is no difficulty in extending these calculations to computing the spin-flip VVCS amplitudes that enter our sum rules. Our results and their mixed level of agreement with EFT predictions are strong incentives to extend these calculations to spin-dependent VVCS amplitudes, allowing lattice QCD predictions to be extensively tested.
@@
Apart from tests of fundamental predictions in the hadronic spin sector, our data will also provide input for calculations of two-photon effects in the hyperfine splitting of the hydrogen (or muonic hydrogen) ground state. The lack of proton spin data at low enough has made this contribution the dominant source of uncertainty in that calculation. The EG4 longitudinal spin data combined with the transverse spin data from the recent JLab Hall A experiment E08-027 [163] will help to remedy this situation [166].
@@ VIII Summary and Conclusions
In this archival paper, we present the details of the Jefferson Lab EG4 experiment and our final results for proton and the deuteron spin structure function data for values as low as 0.012 and 0.017 GeV2, respectively. We include results on the proton and the deuteron , , and their moments, reported previously in Refs. [21, 22], along with details of the simulation and data analysis procedure that led to these results. Our results agree well with previous JLab experiments, but the significantly improved precision of the new data reveal that the world knowledge on low spin structure functions, mostly encapsulated in the parameterization [102], must be updated.

### v4

@@ VII.6.2 Isospin Study of
Our results on the isovector and isoscalar combination of from EG4 are shown in Fig. 26. We also provide results based on EG4 data and E97-110 (3He) data [E97-110:2021mxm]. The total experimental systematic uncertainties on the proton and deuteron/neutron(3He) are combined quadratically, as are the statistical uncertainties. The low- uncertainties are added linearly. The results shown in Fig. 26 agree reasonably well with those from the previous experiments EG1b and EG1b/E94-010 [Deur:2008ej]. However, unlike for , results from the EG4-only and EG4/E97-110 approaches agree only at the higher points. On the other hand, if we combine linearly the total systematic uncertainties of the two experiments, the tension at low vanishes.
@@ VII.7 Discussions
As discussed in previous sections, the success of EFT in describing our results on the moments of the spin structure function remains limited. One reason for this limited success may be coming from the difficulty to fully account for the degree of freedom in the calculation. In fact, early EFT calculations [Bernard:1992nz, Ji:1999pd] did not include explicitly the excitation, which slows the convergence of the EFT perturbation series, or they included it phenomenologically [Bernard:2002bs, Kao:2002cp]. This was thought to not be important for some observables for which the contribution was expected to be suppressed, such as the longitudinal-transverse spin polarizability and the isovector quantities [Burkert:2000qm] or . It therefore came as a surprise that much of the early nucleon spin structure function data [Amarian:2002ar] disagreed with calculations [Bernard:1992nz, Ji:1999pd, Bernard:2002bs, Kao:2002cp], noticeably and . This puzzle prompted refined EFT calculations [Bernard:2012hb, Alarcon:2020icz] and a new experimental program at JLab optimized to cover the EFT domain [CLAS:2017ozc, JeffersonLabE97-110:2019fsc, JeffersonLabHallAg2p:2022qap], including the EG4 experiment reported here.
@@
The latest EFT calculations by Bernard et al. [Bernard:2012hb] and Alarcón et al. [Alarcon:2020icz] both included the but differ in their expansion method for the (pion)- corrections. In EFT, the general expansion parameter is , where GeV and GeV is the chiral symmetry breaking scale. To explicitly account for the , the nucleon- mass difference GeV must be included in the chiral expansion. Bernard et al. [Bernard:2012hb] treated as a small parameter of the same order as . Alarcón et al. [Alarcon:2020icz] used as an intermediate scale so that is the expansion parameter for the - corrections. This is referred to as the “-expansion scheme”. Additionally, Alarcón et al. included a phenomenological form factor, viz not derived within EFT, that suppresses the -dependence of observables at large . This clearly extends the range of applicability of the Alarcón et al. prediction [Alarcon:2020icz] to higher compared to that of Bernard et al. [Bernard:2012hb], albeit by introducing an extra scale factor.
@@
An important outcome that can be drawn from the results shown in this section is that EFT, although successful in many instances, remains challenged by results from dedicated polarized experiments at low . This includes not only EG4, but also other experiments that used different detectors, methods, and spin observables [JeffersonLabE97-110:2019fsc, E97-110:2021mxm, JeffersonLabHallAg2p:2022qap]. To address this problem, further improvements on EFT would be required, although it is very difficult to upgrade the predictions to the next order. Since EFT is the leading QCD-rooted approach to calculating the effects of the strong force at large distances, this difficulty poses a problem in our pursuit of a complete description of nature. Alternatively or complementarily, it would be helpful to have predictions from other non-perturbative approaches to QCD, such as lattice QCD, the Dyson-Schwinger Equations, or the AdS/CFT methods. Recent progress in calculating doubly-virtual Compton scattering (VVCS) amplitudes in lattice QCD now makes moments involved in sum rules within reach of this method, see, e.g., Refs. [Lee:2023lnx, Wang:2023omf] for recent development. At present, only calculations for spin-independent VVCS amplitudes are available; these calculations are difficult because they involve computing 4-point correlation functions, in contrast to usual Lattice QCD calculations that require less computing-extensive 2- and 3-point correlations functions. Only recently, advances in methods and computer power have allowed such a computation. There is no difficulty in extending these calculations to computing the spin-flip VVCS amplitudes that enter our sum rules. Our results and their mixed level of agreement with EFT predictions are strong incentives to extend these calculations to spin-dependent VVCS amplitudes, allowing lattice QCD predictions to be extensively tested.
@@
Apart from tests of fundamental predictions in the hadronic spin sector, our data will also provide input for calculations of two-photon effects in the hyperfine splitting of the hydrogen (or muonic hydrogen) ground state. The lack of proton spin data at low enough has made this contribution the dominant source of uncertainty in that calculation. The EG4 longitudinal spin data combined with the transverse spin data from the recent JLab Hall A experiment E08-027 [JeffersonLabHallAg2p:2022qap] will help to remedy this situation [Ruth:2024bsl].
@@ VIII Summary and Conclusions
In this archival paper, we present the details of the Jefferson Lab EG4 experiment and our final results for proton and the deuteron spin structure function data for values as low as 0.012 and 0.017 GeV2, respectively. We include results on the proton and the deuteron , , and their moments, reported previously in Refs. [CLAS:2017ozc, CLAS:2021apd], along with details of the simulation and data analysis procedure that led to these results. Our results agree well with previous JLab experiments, but the significantly improved precision of the new data reveal that the world knowledge on low spin structure functions, mostly encapsulated in the parameterization [CLAS:2017qga], must be updated.